# Designing an Optimal Sensor Network via Minimizing Information Loss

### v1

@@ 7 Discussion & Conclusions
Key to the scalability of our method is a Gauss-Markov prior on the spatiotemporal process, which recent work has shown has a tractable and efficient variational solution when the inducing points are spatially fixed. Using the sparse formulation given by GPs, we avoid combinatorial optimization of sensor locations, instead solving a relaxed and continuous optimization problem amendable to gradient-based inference.
The use of a Markovian GP has significant advantages as it connects to a rich classical literature on BED, allowing for the incorporation of existing sensors and straightforward extensions to multi-fidelity data. One application of multi-fidelity approaches within our framework is using additional data from new deployments, where scientific data not belonging to the designed sensor network may become available. For example, in climate science applications, measurements are often targeted at a particular area or structure, such as parks in urban areas. One advantage of our proposed method is that it directly provides a predictive GP as an output, which has been well studied as a surrogate model and offers established methods to mix real and simulated data. The most popular approach with GPs in this scenario is to model the error of the physics-based simulator itself with another GP, introduced by kennedy2001bayesian. Since the sum of two GPs is itself a GP, this method retains the convenient predictive structure of our approach.
In our case study, we show that the resulting optimal sensor locations lead to significant advantages over random or quasi-random sensor locations. We further show that, with respect to physics-based model data, the resulting surrogate model predictions are accurate and well-calibrated. This suggests that our method, combined with real observations from the optimized sensor network, may be used as a powerful tool in evaluating the performance of computationally expensive physics-based models in the sense that large disagreements with field observations can be attributed to deficiencies in the physics employed. This is scientifically valuable, as it provides a cheap way to interrogate physics-based models and identify areas where their improvement is necessary.
While the methodological tools of Markovian GPs and variational inference are key to the development of our method, they also introduce technical constraints that are interesting topics for future work. For example, the ST-SVGP work of hamelijnck2021spatio relies on the separability assumption of the kernel, which is unlikely to hold in many applications. One path for future research on this problem is to consider mixture-of-experts that draw from prior knowledge of temporal nonstationarity, e.g., by considering different models for the times in which an urban area is “heating” or “cooling.” In a similar fashion, we may consider introducing more complex spatial kernels in future applications, e.g., through deep kernel learning (wilson2016deep).
@@ Appendix A Further Discussion of Bayesian Optimal Experimental Design
This criterion, known as maximum entropy sampling (MES), notably does not depend on the observed data (shewry1987maximum; sebastiani2000maximum). This is computationally convenient but philosophically puzzling, since the point of a statistical approach in the first place is to provide observation-driven sensor locations rather than geometrically heuristic sensor locations.
@@
shewry1987maximum; sebastiani2000maximum motivate MES as being equivalent to EIG if the entropy $H(\boldsymbol{\mathbf{Y}}_{\text{test}})$ is independent of the design parameters. However, when applied to the sensor design problem, $H(\boldsymbol{\mathbf{Y}}_{\text{test}})$ clearly depends on the sensor locations $\mathbf{X}_{n}$, since $\mathbf{X}_{\text{test}}\triangleq\mathcal{X}\setminus\mathbf{X}_{n}$. Thus, we cannot claim that MES is equivalent to maximizing the EIG.
@@
The classical Bayesian D-optimality (chaloner1995Bayesian) has a complementary form,
@@ Appendix B Illustrating The Incorporation of Existing Sensors
In this appendix, we illustrate the incorporation of existing sensor locations into the MIL framework. In particular, in Fig. 10, we sketch the evolution of sensor locations with the existing locations present. We also include a real example, based on our case study, that shows the effect on learned sensor networks with and without an existing sensor placed in the center of the domain, reinforcing effectiveness.
@@ Appendix C Baselines for Case Study
We compare two popular strategies in spatial OED, also employed in the spatio-temporal setting (mateu2012spatio), namely (a) maximizing the entropy of the distribution of the observations at the design (jin2012entropy), and (b) minimizing the variance of the predictions given the design (heuvelink2012sampling). We refer to the former as MES, and to the latter as IMSE. Certainly, these model-based criteria depend on the chosen model. To keep comparisons fair, we choose the same spatio-temporal GP prior used in MIL (i.e., the ST-SVGP (hamelijnck2021spatio), using the same kernels as in the main manuscript). Furthermore, these two strategies require an initial set of real observations to estimate model hyperparameters and compute a predictive distribution. This is fundamentally different from our approach, where we assume we are in a pre-observation setting and use synthetic data. Hence, for a fair comparison in MES and IMSE, we should use real data both for finding the optimal design and for evaluating how good it is. On the other hand, in MIL, we should use only synthetic data and then evaluate it against real observations. Since we are using a single synthetic dataset, the results in this experiment slightly favor our MIL.
@@ C.1 Maximum Entropy Sampling
In MES, we initialize the model by fitting it to a dataset containing temperatures at $N_{\text{init}}=5$ spatial locations during the first month (June). These initial points were chosen in a space-filling fashion by running the k-means algorithm on the whole spatial grid and selecting the grid points closer to each centroid (jin2017k). Then, the predictive distribution (which can also be seen as a type of conditional prior) over the next month (July) is used to select the design. More specifically, we choose $N_{\text{add}}\in\{1,4,7,10\}$ such that the log-determinant of the posterior covariance matrix at the input design is maximized. The total number of sensors $N_{\text{tot}}=N_{\text{init}}+N_{\text{add}}$ is $N_{\text{tot}}\in\{6,9,12,15\}$.
To ensure computational traceability, we solve the MES in a continuous space. One well-known drawback of MES is its proclivity to push points to the boundary of the design space – to create a motivated design space, we compute the convex hull over all spatial locations and use projected gradient descent (bertsekas1999nonlinear, Ch. 2) for tractability.
Due to the separability assumption on the kernel and the prediction set being a spatiotemporal grid, the full kernel may be symbolically constructed as a Kronecker product, $K=\Sigma_{S}\otimes K_{T}$, for some covariance matrix $\Sigma_{S}$ involving only spatial elements (hamelijnck2021spatio, Eq. (27)). The log-determinant thus decomposes as $\log\det K=N_{T}\log\det\Sigma_{S}+N_{S}\log\det K_{T}$ (minka2000old, Sec. 2); since the second part of this sum does not depend on the spatial design, we may optimize purely with respect to $\log\det\Sigma_{S}$. This allows for computational tractability, as it does not materialize a large covariance matrix or compute its determinant. It further highlights a drawback of the MES approach, namely that it does not capture any temporal information in its spatial designs.
@@ C.2 Integrated Minimum Square Error
In IMSE, we proceed similarly and fit the model to obtain the predictive distribution over the next month (July). The design is chosen such that the trace of the posterior covariance matrix, augmented with “imagined” observations from July at a set of predefined locations is minimum. Notably, because of the structure of GP predictive variances, it does not matter what these imagined observations are, so long as the kernel is fixed. This allows us to compute the IMSE-optimal design exactly with Markovian Gaussian processes. This, in turn, provides $\mathcal{O}(N_{t})$ scaling over the design, which offers a computational advantage over using a conventional GP in spatiotemporal design. This is, to the best of our knowledge, a methodologically novel improvement.
Note that for computing this criterion, we need to choose a fixed “prediction” grid to “integrate” over, conditional on the chosen design. We chose the entire spatial grid as the prediction grid (314 points). This aligns with the evaluation criteria (both RMSE and NPLL), which are computed over the same grid.
@@ Appendix D Preliminary Experiments With Deep Kernels
In Section 6, we included a case study using a separable kernel, with a mixture Hida-Matérn kernel as the temporal sub-kernel and a Matérn kernel for the spatial sub-kernel. While separability and a stationary temporal kernel are key assumptions we make for convenient variational algorithms, it was noted that the spatial kernel need not be stationary. Indeed, the spatial kernel can be an arbitrary positive semi-definite kernel, e.g., as learned via deep kernel learning (wilson2016deep).
@@
We repeat the experiment in Section 6, training using a Matérn-3/2 kernel and deep Matérn-3/2 kernel for $250$ epochs, using a reducing learning rate schedule for the deep Matérn-3/2 kernel. We obtain visually distinct optimal experimental designs (Fig. 12), with somewhat degraded performance (Fig. 12). We visualize the spatial non-stationarity of the resulting kernel in Fig. 13. In our experiments, we find that deep kernels are, somewhat unsurprisingly, more sensitive to hyperparameters and less stable than their stationary counterparts.
@@
Figure 13: A comparison of the learned Matérn-3/2 and Deep Matérn-3/2 spatial kernels. The Matérn-3/2 exhibits some mild non-stationarity, while the Matérn-3/2 kernel does not.
@@ Appendix E Experiments with “Simulator Error”
To further investigate the robustness of the MIL selection criterion under simulator misspecification, we conducted a controlled numerical study. Specifically, we construct a toy example based on our Phoenix, Arizona case study by adding GP-based simulator error, i.e., $f_{\text{obs}}(\boldsymbol{\mathbf{x}},t)=f_{\text{true}}(\boldsymbol{\mathbf{x}},t)+f_{\text{err}}(\boldsymbol{\mathbf{x}},t)$. This is in line with classical assumptions of simulator error (kennedy2001bayesian), with noise in our case being drawn from a separable Matérn-3/2 kernel,
@@
$=\sigma_{\text{noise}}^{2}\kappa_{\text{Mat-3/2}}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{x}}^{\prime};\ell_{x})\kappa_{\text{Mat-3/2}}(t,t^{\prime};\ell_{t}).$ (37)
To evaluate the impact of misspecification, we compared the design points selected by MIL across the different “simulators.” To better measure how simulator error directly influences performance, we chose the training set to be exactly the evaluation set, with the added GP noise. The noise uses spatial lengthscales – in normalized units – of $\ell_{s}\in\{0.1,1.0\}$, representing localized and spatially diffuse noise, respectively, and temporal noise in $\ell_{t}\in\{1.0,36.0\}.$ The training set comprises the first 7 days of data from June 2013. We test combinations of small and large noise variances, with localized and globalized spatial and temporal lengthscales. Each experiment introduces $9$ sensors.
We extract a set of “golden” hyperparameters by first fitting an ST-SVGP to the ground-truth data to isolate predictive error stemming from hyperparameter learning and predictive error from inferior sensor networks. We repeat for 10 replications, which we note provide legitimately different noise characteristics each time.
Results, in terms of RMSE and NPLL, are shown in Fig. 14. The results support our theoretical arguments, i.e., that moderate simulator error results in acceptable error rates, still comparable to baseline methods, and that localized noise and high signal-to-noise ratios are both crucial. We additionally visualize the sensor locations learned in each noise setting for a specific realization in Fig. 15. We can easily verify that the noise variance has legitimate effects on the resulting sensor locations.

### v2

@@ 7 Discussion & Conclusions
Key to the scalability of our method is a Gauss-Markov prior for the spatiotemporal process, for which recent work has shown that efficient and tractable variational inference is possible when inducing points are spatially fixed. Using the sparse formulation given by GPs, we avoid the combinatorial optimization of sensor locations and instead solve a relaxed, continuous optimization problem amenable to gradient-based inference.
The use of a Markovian GP has significant advantages as it connects to a rich classical literature on BED, allowing for the incorporation of existing sensors and straightforward extensions to multi-fidelity data. One application of multi-fidelity approaches within our framework is using additional data from new deployments, where scientific data not belonging to the designed sensor network may become available. One advantage of our proposed method is that it directly provides a GP predictive model as a byproduct, which has been well studied as a surrogate model and offers established methods to mix real and simulated data. The most popular approach with GPs in this scenario is to model the error of the physics-based simulator itself with another GP, introduced by Kennedy and O’Hagan (2001). Since the sum of two GPs is itself a GP, this method retains the convenient predictive structure of our approach.
In our case study, we show that the resulting optimal sensor locations lead to significant advantages over existing approaches. We further show that, with respect to physics-based model data, the resulting surrogate model predictions are accurate and well-calibrated. This suggests that our method, combined with real observations from the optimized sensor network, may be used as a powerful tool for evaluating the performance of computationally expensive physics-based models, in the sense that large disagreements with field observations can be attributed to deficiencies in the physics employed. This is scientifically valuable, as it provides a cost-effective way to interrogate physics-based models and identify areas where their improvement is necessary.
While the methodological tools of Markovian GPs and variational inference are key to the development of our method, they also introduce technical constraints that are interesting topics for future work. For example, the ST-SVGP work of Hamelijnck et al. (2021) relies on the separability assumption of the kernel, which is unlikely to hold in many applications. One path for future research on this problem is to consider mixture-of-experts that draw from prior knowledge of temporal nonstationarity, e.g., by considering different models for the times when an urban area is “heating” or “cooling.” In a similar fashion, we may consider introducing more complex spatial kernels in future applications, e.g., through deep kernel learning (Wilson et al., 2016).
@@ Appendix A Further Discussion of Bayesian Optimal Experimental Design
This criterion, known as maximum entropy sampling (MES), notably does not depend on the observed data explicitly (Shewry and Wynn, 1987; Sebastiani and Wynn, 2000). This is computationally convenient but philosophically puzzling, since the point of a statistical approach in the first place is to provide observation-driven sensor locations rather than geometrically heuristic sensor locations.
@@
Shewry and Wynn (1987); Sebastiani and Wynn (2000) motivate MES as being equivalent to EIG if the entropy $H(\boldsymbol{\mathbf{Y}}_{\text{test}})$ is independent of the design parameters. However, when applied to the sensor design problem, $H(\boldsymbol{\mathbf{Y}}_{\text{test}})$ clearly depends on the sensor locations $\mathbf{X}_{n}$, since $\mathbf{X}_{\text{test}}\triangleq\mathcal{X}\setminus\mathbf{X}_{n}$. Thus, we cannot claim that MES is equivalent to maximizing the EIG.
@@
The classical Bayesian D-optimality (Chaloner and Verdinelli, 1995) admits a prediction-space analogue in the GP setting,
@@ Appendix B Illustrating The Incorporation of Existing Sensors
In this appendix, we illustrate the incorporation of existing sensor locations into the MIL framework. In particular, in Fig. 10, we sketch the evolution of sensor locations with the existing locations present. We also include a real example based on our case study that shows the effect on learned sensor networks with and without an existing sensor placed in the center of the domain, reinforcing effectiveness.
@@ Appendix C Baselines for Case Study
We compare two popular strategies in spatial OED, also employed in the spatio-temporal setting (Mateu and Müller, 2012), namely (a) maximizing the entropy of the distribution of the observations at the design (Jin et al., 2012), and (b) minimizing the variance of the predictions given the design (Heuvelink et al., 2012). We refer to the former as MES and to the latter as IMSE. Certainly, these model-based criteria depend on the chosen model. To keep comparisons fair, we choose the same spatio-temporal GP prior used in MIL (i.e., the ST-SVGP (Hamelijnck et al., 2021), using the same kernels as in the main manuscript). Furthermore, these two strategies require an initial set of real observations to estimate model hyperparameters and compute a predictive distribution. This is fundamentally different from our approach, where we assume we are in a pre-observation setting and use synthetic data. Hence, for a fair comparison in MES and IMSE, we should use real data both to find the optimal design and to evaluate its effectiveness. On the other hand, in MIL, we should use only synthetic data and then evaluate it against real observations. Since we are using a single synthetic dataset, the results of this experiment slightly favor our MIL.
@@ C.1 Maximum Entropy Sampling
In MES, we initialize the model by fitting it to a dataset containing temperatures at $N_{\text{init}}=5$ spatial locations during the first month (June). These initial points were chosen in a space-filling fashion by running the k-means algorithm on the entire spatial grid and selecting the grid points closer to each centroid (Jin and Han, 2017). Then, the predictive distribution (which can also be seen as a type of conditional prior) over the next month (July) is used to select the design. More specifically, we choose $N_{\text{add}}\in\{1,4,7,10\}$ such that the log-determinant of the posterior covariance matrix at the input design is maximized. The total number of sensors $N_{\text{tot}}=N_{\text{init}}+N_{\text{add}}$ is $N_{\text{tot}}\in\{6,9,12,15\}$.
To ensure computational traceability, we solve the MES in a continuous space. One well-known drawback of MES is its proclivity to push points to the boundary of the design space – to create a motivated design space, we compute the convex hull over all spatial locations and use projected gradient descent (Bertsekas, 1999, Ch. 2) for tractability.
Due to the separability assumption on the kernel and the prediction set being a spatiotemporal grid, the full kernel may be symbolically constructed as a Kronecker product, $K=\Sigma_{S}\otimes K_{T}$, for some covariance matrix $\Sigma_{S}$ involving only spatial elements (Hamelijnck et al., 2021, Eq. (27)). The log-determinant thus decomposes as $\log\det K=N_{T}\log\det\Sigma_{S}+N_{S}\log\det K_{T}$ (Minka, 2000, Sec. 2); since the second part of this sum does not depend on the spatial design, we may optimize purely with respect to $\log\det\Sigma_{S}$. This allows for computational tractability, as it does not materialize a large covariance matrix or compute its determinant. It further highlights a drawback of the MES approach, namely that it does not capture any temporal information in its spatial designs.
@@ C.2 Integrated Minimum Square Error
In IMSE, we proceed similarly and fit the model to obtain the predictive distribution over the next month (July). The design is chosen such that the trace of the posterior covariance matrix, augmented with “imagined” observations from July at a set of predefined locations, is minimized. Notably, because of the structure of GP predictive variances, it does not matter what these imagined observations are, so long as the kernel is fixed. This allows us to compute the IMSE-optimal design exactly with Markovian Gaussian processes. This, in turn, provides $\mathcal{O}(N_{t})$ scaling over the design, which offers a computational advantage over using a conventional GP in spatiotemporal design. This is, to the best of our knowledge, a methodologically novel improvement.
Note that for computing this criterion, we need to choose a fixed prediction grid to sum over, conditional on the chosen design. We chose the entire spatial grid as the prediction grid (314 points). This aligns with the evaluation criteria (both RMSE and NPLL), which are computed over the same grid.
@@ Appendix D Preliminary Experiments With Deep Kernels
In Section 6, we included a case study using a separable kernel, with a mixture Hida-Matérn kernel as the temporal sub-kernel and a Matérn kernel for the spatial sub-kernel. While separability and a stationary temporal kernel are key assumptions we make for convenient variational algorithms, it was noted that the spatial kernel need not be stationary. Indeed, the spatial kernel can be an arbitrary positive semi-definite kernel, e.g., as learned via deep kernel learning (Wilson et al., 2016).
@@
We repeat the experiment in Section 6, training using a Matérn-3/2 kernel and a deep Matérn-3/2 kernel for $250$ epochs, using a reducing learning rate schedule for the deep Matérn-3/2 kernel. We obtain visually distinct optimal experimental designs (Fig. 12), with slightly degraded performance (Fig. 12). We visualize the spatial non-stationarity of the resulting kernel in Fig. 13. In our experiments, we find that deep kernels are, somewhat unsurprisingly, more sensitive to hyperparameters and less stable than their stationary counterparts.
@@
Figure 13: A comparison of the learned Matérn-3/2 and Deep Matérn-3/2 spatial kernels. The Deep Matérn-3/2 exhibits some mild non-stationarity, while the Matérn-3/2 kernel does not.
@@ Appendix E Experiments with “Simulator Error”
To further investigate the robustness of the MIL selection criterion under simulator misspecification, we conducted a controlled numerical study. Specifically, we constructed a toy example based on our Phoenix, Arizona case study by introducing GP-based simulator error, i.e., $f_{\text{obs}}(\boldsymbol{\mathbf{x}},t)=f_{\text{true}}(\boldsymbol{\mathbf{x}},t)+f_{\text{err}}(\boldsymbol{\mathbf{x}},t)$. This is in line with classical assumptions of simulator error (Kennedy and O’Hagan, 2001), with noise in our case being drawn from a separable Matérn-3/2 kernel,
@@
$=\sigma_{\text{noise}}^{2}\kappa_{\text{Mat-3/2}}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{x}}^{\prime};\ell_{s})\kappa_{\text{Mat-3/2}}(t,t^{\prime};\ell_{t}).$ (36)
To evaluate the impact of misspecification, we compared the design points selected by MIL across different simulator realizations. To better measure how simulator error directly influences performance, we chose the training set to be exactly the evaluation set, with the added GP noise. The noise uses spatial lengthscales – in normalized units – of $\ell_{s}\in\{0.1,1.0\}$, representing localized and spatially diffuse noise, respectively, and temporal noise in $\ell_{t}\in\{1.0,36.0\}.$ The training set consists of the first 7 days of June 2013. We evaluate combinations of small and large noise variances with localized and globalized spatial and temporal lengthscales. Each experiment deploys $9$ sensors.
We extract a set of “golden” hyperparameters by first fitting an ST-SVGP to the ground-truth data to isolate predictive error stemming from hyperparameter learning and predictive error from inferior sensor networks. We repeat this process for 10 replications, which we note provide legitimately different noise characteristics each time.
Results, in terms of RMSE and NLPD, are shown in Fig. 14. The results support our theoretical arguments, i.e., that moderate simulator error results in acceptable error rates, still comparable to baseline methods, and that localized noise and high signal-to-noise ratios are both crucial. We additionally visualize the sensor locations learned in each noise setting for a specific realization in Fig. 15. We can easily verify that the noise variance has legitimate effects on the resulting sensor locations.